Determine when matthew first ranked his intensity at 7.5. f(x)=5sin 3.14/2(n-2)+5


1
Expert's answer
2021-06-21T15:40:17-0400

"f(x)=7.5"

"5sin[\frac{3.14}{2}(n-2)]+5=7.5"

Subtract 5 both sides:

"5sin[\frac{3.14}{2}(n-2)]+5-5=7.5-5"

"5sin[\frac{3.14}{2}(n-2)]=2.5"

Divide by 5 both sides

"sin[\frac{3.14}{2}(n-2)]=2.5÷5"

"sin[\frac{3.14}{2}(n-2)]=\frac{1}{2}"

"\frac{\pi}{2}(n-2)=\frac{π}{6}\\ wherebyπ=3.14"

"n-2=\frac{1}{3}"

Add 2 both sides

"n=\frac{1}{3}+2\\n=\frac{7}{3}"